### 28 Apr 2026 v2

# Improving LLM Predictions via Inter-Layer Structural Encoders

Abstract
The standard practice in Large Language Models (LLMs) is to base predictions on final-layer representations. However, intermediate layers encode complementary task-relevant signals, and the optimal layer is task-dependent, making single-layer usage inherently suboptimal. In this work, we introduce Inter-Layer Structural Encoders (ILSE), a powerful and parameter-efficient post-training framework that learns to aggregate representations from all layers of a frozen LLM through structured inter-layer interactions. Central to ILSE is the Cayley-Encoder, a mathematically grounded module based on expander Cayley graphs that enables efficient and effective inter-layer information propagation. We evaluate ILSE on 13 classification and semantic similarity tasks across 9 pre-trained LLMs ranging from 14M to 8B parameters. ILSE consistently outperforms strong baselines, achieving up to 44% improvements in accuracy and 25% in similarity, while introducing at most 0.1% additional parameters relative to the base LLM size. Furthermore, ILSE is highly data-efficient in few-shot regimes and enables small LLMs to match or exceed the performance of substantially larger models. Notably, it also outperforms LoRA-based fine-tuning despite operating on frozen representations.
1 Introduction
Large Language Models (LLMs) have transformed natural language processing through their ability to learn rich, transferable representations (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020). In practice, downstream predictions are typically based on the final-layer representations, under the assumption that deeper layers contain the most task-relevant information. However, recent work challenges this view, showing that intermediate layers encode complementary signals and can outperform the final layer depending on the task (Wallat et al., 2020; Fan et al., 2024; Skean et al., 2025).
A growing body of work has demonstrated that linguistic information is distributed across layers: lower layers capture surface-level features, middle layers encode syntactic structure, and higher layers focus on semantics (Tenney et al., 2019; Jawahar et al., 2019; Clark et al., 2019). Consequently, the optimal representation is inherently task-dependent, making reliance on a single layer suboptimal. This raises a fundamental question: how can representations from all layers be combined effectively? Several methods have proposed aggregating representations from all layers to improve downstream performance (Peters et al., 2018; ElNokrashy et al., 2024; Pagliardini et al., 2024; Xiao et al., 2025). However, these approaches either rely on heuristic or constrained aggregation mechanisms, or require modifying the transformer architecture and retraining it from scratch, incurring substantial computational cost.
In this work, we introduce Inter-Layer Structural Encoders (ILSE), a parameter-efficient post-training framework that aggregates representations from all layers of a frozen LLM into a single, improved representation. ILSE treats layer representations as elements of a structured object and learns their interactions using either DeepSets (Zaheer et al., 2017) or Graph Neural Networks (GNNs) (Gilmer et al., 2017).
In contrast to prior approaches, ILSE adopts a principled and mathematically grounded formulation rooted in recent advances in graph learning. In particular, it avoids overfitting to the chosen structure by leveraging regular graphs (Bechler-Speicher et al., 2024), while enabling efficient and bottleneck-free inter-layer communication via expander Cayley graphs (Alon and Yahav, 2021; Wilson et al., 2025). Concretely, ILSE employs structured inter-layer aggregation based on three regimes: (1) Cayley graphs over the Special Linear Group $SL(2,\mathbb{Z}_{n})$, which combine sparsity with strong connectivity guarantees, (2) fully connected graphs, and (3) sets.
We evaluate ILSE on 13 classification and semantic text similarity (STS) tasks across 9 pre-trained LLMs ranging from 14M to 8B parameters. ILSE consistently outperforms existing approaches by large margins, while remaining highly parameter-efficient and data-efficient.
Our main contributions are as follows.
  1. 1.
    We propose ILSE, a principled and mathematically grounded, parameter-efficient post-training framework for structured aggregation of representations across all layers of a frozen LLM.
  2. 2.
    In an extensive evaluation of ILSE across 9 pre-trained LLMs and 13 classification and semantic similarity tasks, we show consistent improvements over strong baselines, achieving up to 44% improvements in accuracy and 25% in similarity, while introducing at most 0.1% additional parameters relative to the base LLM size.
  3. 3.
    We perform a few-shot evaluation with increasing amount of samples, we establish the data efficiency of ILSE, achieving strong performance with as few as 32 samples per label and outperforming full-data baselines in low-resource regimes.
  4. 4.
    We perform an evaluation of ILSE over a wide range of LLM sizes, ranging from 14 millionth 2.8 billion parameters, showing that it enables small LLMs to match or exceed the performance of substantially larger ones.
  5. 5.
    We show that ILSE outperforms LoRA-based fine-tuning.
2 Related Work
2.1 Aggregated Representations in LLMs
Research into transformer layers has revealed that different layers capture different linguistic features. Jawahar et al. (2019) showed that BERT organizes linguistic knowledge hierarchically: surface features in lower layers, syntactic features in middle layers, and semantic features in upper layers. Tenney et al. (2019) found that syntactic information concentrates in specific layers while semantic information spreads across multiple layers. Wallat et al. (2020) further showed that often intermediate layers capture factual knowledge better than the last layer, while Skean et al. (2025) demonstrated that intermediate layer representations consistently outperform final layer representations across downstream tasks. They propose an unsupervised approach for optimal layer selection for downstream tasks.
Several approaches leverage information from multiple layers. Peters et al. (2018) introduced task-specific scalar weights for each layer representation with ELMo, applied over BiLSTM layers, effectively treating them as a linear combination of independent features. More recently, ElNokrashy et al. (2024) extended this to transformers with Depth-Wise Attention (DWAtt). DWAtt uses a query derived from the last-layer hidden state to attend over depth, with learned positional keys for layers and MLP-transformed values from each intermediate layer representation. Both approaches suffer from limitations: ELMo restricts aggregation to a simple linear mixture, missing non-linear interactions between layers. While DWAtt models non-linear interactions, it explicitly constructs the attention query from the final layer representation, utilizing previous layers only as keys and values. This approach focuses more on the final layer rather than equal treatment for all layers.
Other work modifies the transformer architecture itself. Pagliardini et al. (2024) introduce weighted averaging modules after each transformer block, while Xiao et al. (2025) propose dynamic, position-dependent layer mixing. However, these architectural changes require training the transformer from scratch.
Our work addresses these limitations: we enable structured inter-layer communication within a frozen transformer, without modifying the architecture (Figure  1).
Figure 1: An overview of ILSE with Clayey-Encoder. Layer representations are mapped into nodes in a Cayley Graph, which is then fed into a GNN to learn the final inter-layer representation.
2.2 Graph Neural Networks and Cayley Graphs
Graph Neural Networks (GNNs) are a family of neural networks designed to learn representations over graph-structured data. GNNs operate by iteratively propagating information between nodes through a process called message passing neural network (MPNN)  (Gilmer et al., 2017). Each MPNN layer consists of three operations: (1) each node computes messages to send to its neighbors, (2) each node aggregates the messages it receives, and (3) each node updates its representation by combining the aggregated messages with its previous state.
Formally, consider a graph $G=(V,E)$ with nodes $V$ and edges $E$. Each node $v$ is associated with an initial representation $h_{v}^{(0)}\in\mathbb{R}^{d}$, typically derived from input features. At layer $k$, the representation of node $v$ is updated as follows:
$m_{v}^{(k)}$ $=\sum_{u\in\mathcal{N}(v)}M^{(k)}\left(h_{v}^{(k-1)},h_{u}^{(k-1)}\right)$ (1)
$h_{v}^{(k)}$ $=U^{(k)}\left(h_{v}^{(k-1)},m_{v}^{(k)}\right)$ (2)
where $m_{v}^{(k)}$ is the aggregated message received by node $v$ at layer $k$, $\mathcal{N}(v)$ denotes the set of neighbors of node $v$, $M^{(k)}$ is the message function that computes the contribution from each neighbor, and $U^{(k)}$ is the update function that combines the node’s previous representation with its aggregated messages. Both $M^{(k)}$ and $U^{(k)}$ are learnable functions. After $K$ layers of message passing, the final node representations $h_{v}^{(K)}$ can be used directly for node-level tasks. To form a graph-level representations the final node representations aggregated using a permutation-invariant function, possibly followed by another learned function usually denoted as a readout.
Many GNN variants have been proposed, differing mostly in how they aggregate information. For example, Graph Convolutional Networks (GCNs) (Kipf and Welling, 2017) use a normalized mean aggregation, where messages are weighted by node degrees. Graph Isomorphism Networks (GINs) (Xu et al., 2019) aggregate messages using summation followed by a multilayer perceptron (MLP) (Hinton, 2007), and were proven to be maximally expressive within the MPNN family.
A limitation of standard GNNs is over-squashing (Alon and Yahav, 2021). As the number of layers grows, information propagating through the graph is compressed exponentially into fixed-size node representations, creating bottlenecks that reduce the expressivity of GNNs and hinder long-range communication between nodes (Alon and Yahav, 2021; Di Giovanni et al., 2023).
Deac et al. (2022) introduced Expander Graph Propagation (EGP) to improve information flow in GNNs. They address over-squashing by leveraging Cayley graphs. Cayley graphs are sparse yet highly connected graph structures with logarithmic diameter, allowing any two nodes to communicate in $O(\log|V|)$ hops. These properties make them effective for propagating information without bottlenecks. EGP constructs a Cayley graph and alternates message passing between the original input graph and the Cayley graph. Cayley graphs come in fixed sizes determined by the underlying group, so EGP truncates them to match the input graph size. However, Wilson et al. (2025) showed that truncation can reintroduce bottlenecks. They proposed Cayley Graph Propagation (CGP), which instead pads the input graph with virtual nodes to preserve the complete Cayley structure, thereby maintaining the connectivity guarantees and avoiding the bottlenecks introduced by truncation. The virtual nodes are nodes used only to propagate information of the graph, but are not used for the final aggregation.
3 Inter-Layer Structural Encoders
In this section, we introduce ILSE, an effective and mathematically grounded approach for combining inter-layer representations into one enhanced representation.
Preliminaries
We consider LLMs with $L$ transformer layers and an input sequence of $T$ tokens. Each layer $\ell\in\{1,\ldots,L\}$ produces a $d$-dimensional hidden representation for each token $\mathbf{H}_{\ell}\in\mathbb{R}^{T\times d}$. We form a layer representation $\mathbf{z}_{\ell},\ell\in\{1,\ldots,L\}$ by averaging the hidden token representation of each layer
$\mathbf{z}_{\ell}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{H}_{\ell}[t,:]\in\mathbb{R}^{d}$
A Cayley graph is a graph that describes the abstract structure of a mathematical group. Here we focus on the special linear group $SL(2,\mathbb{Z}_{n})$, with elements that are defined as $2\times 2$ matrices with integer entries modulo $n$ and unit determinant. Cayley graphs over $SL(2,\mathbb{Z}_{n})$ are 4-regular, and have logarithmic diameter relative to the number of nodes, enabling efficient global communication between nodes when applying message-passing on them. The number of nodes in the graph $|V_{n}|$ is determined by the formula:
$|V_{n}|=n^{3}\prod_{\text{prime }p|n}\left(1-\frac{1}{p^{2}}\right)$ (3)
ILSE introduces three inter-layer structural encoders:
  • Cayley-Encoder
  • Set-Encoder
  • Fully-Connected (FC) Encoder
Cayley-Encoder Cayley-Encoder maps layers to nodes of a Cayley Graph over $SL(2,\mathbb{Z}_{n})$, and then applies a GNN. To maintain the graph’s symmetry and expansion properties, we choose the smallest $n$ such that $|V_{n}|\geq L$. We map the $L$ layer representations randomly to $L$ nodes of the Cayley graph. When $|V_{n}|>L$, the remaining $|V_{n}|-L$ nodes are initialized as virtual nodes with zero vectors as their representations. We then apply a GNN to learn a final representation over the Cayley graph. Following (Wilson et al., 2025), we use only the representations from the non-virtual nodes for the final hidden representation used for the prediction task.
Set-Encoder Set-Encoder uses each layer representation as one set element, and then applies a DeepSet Zaheer et al. (2017) to learn the final representation. A DeepSet is a neural architecture for learning permutation-invariant functions over sets. A DeepSet first applies a shared Multi Layer Perceptron (Hinton, 2007) $\phi$ to each element independently, then aggregates the result with a permutation-invariant pooling PI-Pool, and finally applies another transformation $\rho$ which is an MLP that produces the task-specific representation:
$f(Z)=\rho\left(\text{PI-Pool}_{\ell=1}^{L}\phi(\mathbf{z}_{\ell})\right)$
FC-Encoder FC-Encoder constructs a fully connected graph over the layers, where every layer is connected to every other layer, and then applies a GNN to learn the final representation. The fully-connected graph enables direct communication between any pair of layers. This structure is dense, with $\binom{L}{2}$ undirected edges.
In the next section, we demonstrate the effectiveness of ILSE through extensive evaluation.
4 Experiments
In this section we evaluate ILSE on 5 classification and 8 STS tasks from MTEB (Muennighoff et al., 2023). Our experiments include:
  • Full-data training where we train on all the training data available for the task.
  • An evaluation of zero-shot transfer capabilities on STS tasks.
  • A fine-grained performance evaluate as a function of the LLM size, ranging between 14 million and 2.8 billion parameters.
  • Few-shot training with 1110241024 samples per label.
  • A comparison with LoRA fine-tuning.
  • An evaluation of Cayley-Encoders on token-level instead of layer-level.
4.1 Tasks
We evaluate on two types of tasks from MTEB benchmark (Muennighoff et al., 2023): classification and Semantic Text Similarity (STS).
For the classification tasks, we used Banking77, which is a dataset of online banking queries annotated with 77 intent classes. We also used Emotion, a dataset of English Twitter messages labeled with six emotion categories. In addition, we included MTOP Domain, which focuses on predicting the domain of task-oriented dialogue inputs, and MTOP Intent, which involves predicting the user intent for such inputs. Finally, we used Poem Sentiment, a dataset consisting of poem verses annotated for four sentiment classes(Casanueva et al., 2020; Saravia et al., 2018; Li et al., 2021; Sheng and Uthus, 2020; Enevoldsen et al., 2025). We used the training split for each task to train the model and then evaluated on the test split.
For STS tasks, we used STSBenchmark (May, 2021) training split and evaluate on its test split. We then use the same trained models for zero-shot evaluation on all other English STS tasks in MTEB: STS12, STS13, STS14, STS15, STS16, BIOSSES, and SICK-R (1; E. Agirre, M. Diab, D. Cer, and A. Gonzalez-Agirre (2012); E. Agirre, D. M. Cer, M. T. Diab, A. Gonzalez-Agirre, and W. Guo (2013); A. Bandhakavi, N. Wiratunga, D. P, and S. Massie (2014); E. Biçici (2015); P. Nakov, A. Ritter, S. Rosenthal, F. Sebastiani, and V. Stoyanov (2016); M. Marelli, S. Menini, M. Baroni, L. Bentivogli, R. Bernardi, and R. Zamparelli (2014)).
4.2 Setup
To evaluate the effectiveness of ILSE, we compare it against a diverse set of baselines:
  • Last-Layer: The last-layer representation, as done in (Skean et al., 2025).
  • Best-Layer : The best-performing layer, selected by evaluating each layer independently.
  • Weighted: A learned weighted sum of the layer-representations, similarly to (Peters et al., 2018).
  • MLP Last-Layer: A trained MLP over the last-layer representations.
  • MLP Best-Layer An MLP trained on the selected layer from Best-Layer.
  • DWAtt: A projection to 256 dimensions followed by DWAtt (ElNokrashy et al., 2024). The projection is intended to reduce the large parameter count of applying DWAtt directly to the layer-representations.
Table 1 shows the number of learned parameters learned on top of the frozen LLM for each method. For the Last-Layer and Best-Layer methods, there are no added learned parameters. Our code and experimental configurations are publicly available. [^1] Code: https://github.com/tomulanovski/ILSE
We evaluate ILSE with three LLM families: Pythia-410M (25 layers, 1024-dim) (Biderman et al., 2023), Gemma2-2B (27 layers, 2304-dim) (Team, 2024), and Llama3-8B (33 layers, 4096-dim) (AI@Meta, 2024). All base models remain frozen during training and we only train ILSE and the baselines on top of the frozen representations. For each layer, we use the layer representation obtained by mean-pooling over all token representations. These pooled representations serve as the input to all methods and baselines evaluated in this work.
For classification tasks, we train the encoders jointly with a linear classifier head using cross-entropy loss. The base LLM remains frozen throughout training. For STS tasks, we encode each sentence in a pair through the frozen LLM and our encoders, compute the cosine similarity between the resulting representations, and minimize Mean Squared Error (MSE) loss with respect to the ground-truth similarity score. We use the Adam optimizer (Kingma, 2014) with learning rate and weight decay selected via Optuna (Akiba et al., 2019) hyperparameter optimization on the validation set. For FC-Encoder and Cayley-Encoder, we evaluate two GNN architectures: GIN and GCN. We test different parameters values. For the number of MPNN layers we test 1-2, for dropout we test values 0.0 - 0.3, for learning rate and weight decay we test $10^{-4}$ to $10^{-3}$. For the number of layers in the MLP used inside the GIN aggregation we test 0-2. We keep hidden dimension fixed to 256, and batch size to 64. We choose the best parameters set based on the best validation accuracy and train the final model. For the test evaluation we extract the trained encoder architecture without the classifier head for classification and use it for encoding as part of MTEB evaluation - we report the final accuracy and similarity scores on the test set.
Table 1: Trainable parameters added by each method. All base model parameters remain frozen. Percentages show overhead relative to base model size.
Method Pythia-410M Gemma2-2B Llama3-8B
Frozen Parameters 410M 2B 8B
Weighted 25 27 33
MLP 394K (0.10%) 721K (0.036%) 1.18M (0.015%)
Set-Encoder 394K (0.10%) 721K (0.036%) 1.18M (0.015%)
FC-Encoder 395K (0.10%) 722K (0.036%) 1.18M (0.015%)
Cayley-Encoder 395K (0.10%) 722K (0.036%) 1.18M (0.015%)
DWAtt 2.0M (0.49%) 2.46M (0.12%) 3.3M (0.04%)
4.3 Results
Tables 2 and 4 present our main results. ILSE achieves the best performance in all 15 classification configurations, all 3 supervised STS configurations, and 17 out of 21 zero-shot STS transfer settings.
For classification, ILSE achieves average gains of +30% over Last-Layer baseline and +25% over the Best-Layer baseline, which itself requires evaluating every layer independently. We observe particularly large margins on the Emotion task, where ILSE improves up to +44% and +40% over Last-Layer and Best-Layer respectively on average across all three LLMs. Similar patterns emerge across Banking77 and the MTOP tasks, where ILSE consistently reaches 90–99% accuracy compared to 60–85% for Last-Layer and Best-Layer baselines. For STS tasks, Cayley-Encoder and FC-Encoder improves upon Last-Layer and Best-Layer in 7 out of 8 tasks with Cayley-Encoder achieving the highest scores on average.
[…]
Table 2: Performance comparison of ILSE and baselines across 3 different LLMs on 5 classification tasks. For each task and LLM, we highlight in Bold the best performing setting, in blue the 2nd best and in red the 3rd best. In all 15 cases ILSE gets the highest score. For Pythia-410m FC-Encoder is the most dominant method while in Gemma2-2B and Llama3-8B it’s Cayley-Encoder. In all cases but one ILSE also achieves the best second and third scores.
Base Model Section Method Banking77 Emotion MTOPDomain MTOPIntent PoemSentiment
Pythia-410m Baselines Last Layer 61.17 33.48 80.88 66.97 42.40
Best Single Layer 66.67 35.02 83.78 71.18 45.67
MLP Last Layer 83.84 33.99 96.97 83.75 75.00
MLP Best Layer 41.93 25.41 87.25 70.79 53.94
Weighted 58.50 28.26 79.79 61.68 42.60
DWATT 83.23 58.60 98.03 91.66 70.87
ILSE Set-Encoder 84.23 47.89 97.59 92.21 73.37
FC-Encoder (GIN) 90.10 73.36 98.68 94.32 69.90
FC-Encoder (GCN) 90.65 75.61 98.65 95.04 75.77
Cayley-Encoder (GIN) 89.12 73.83 98.77 94.72 70.87
Cayley-Encoder (GCN) 89.43 66.40 98.67 94.19 69.13
Gemma2-2B Baselines Last Layer 62.16 26.89 80.79 68.02 35.67
Best Single Layer 72.31 31.27 87.09 75.36 42.79
MLP Last Layer 87.47 59.05 98.37 92.73 71.63
MLP Best Layer 59.85 30.26 84.24 73.69 46.83
Weighted 65.40 29.94 85.17 73.20 39.62
DWATT 88.82 61.22 98.39 90.93 67.79
ILSE Set-Encoder 90.03 78.77 98.92 94.37 78.56
FC-Encoder (GIN) 92.39 79.55 99.07 94.47 74.62
FC-Encoder (GCN) 91.71 79.90 98.87 95.34 77.12
Cayley-Encoder (GIN) 92.58 68.38 98.88 96.43 76.44
Cayley-Encoder (GCN) 91.86 69.60 99.16 95.97 83.27
Llama3-8B Baselines Last Layer 68.25 34.23 84.42 73.39 40.96
Best Single Layer 71.93 38.42 89.01 78.17 47.02
MLP Last Layer 86.70 67.67 98.58 92.09 75.00
MLP Best Layer 49.59 21.17 47.64 60.04 35.67
Weighted 66.63 27.94 83.85 71.78 37.79
DWATT 90.04 66.55 97.97 92.41 75.00
ILSE Set-Encoder 87.62 71.04 98.77 95.43 77.02
FC-Encoder (GIN) 92.10 71.64 98.77 95.65 75.96
FC-Encoder (GCN) 92.38 71.03 98.99 96.19 77.98
Cayley-Encoder (GIN) 92.46 71.58 98.98 96.46 76.54
Cayley-Encoder (GCN) 92.85 73.43 99.03 95.90 79.04
Among ILSE encoders, we observe that explicit graph structures (FC-Encoder and Cayley-Encoder) consistently outperform Set-Encoder, indicating that inter-layer connectivity improves aggregation. Cayley-Encoder proves particularly effective for STS tasks, achieving the best score in 17 out of 24 STS tasks configurations, while FC-Encoder shows comparable performance on the classification tasks. This suggests that the sparse, regular connectivity of Cayley graphs is particularly beneficial for semantic similarity. The choice between GIN and GCN aggregation shows task-dependent variation, with GIN generally favoring STS tasks and GCN showing slight advantages on classification.
Since Cayley-Encoder maps layer representations randomly to nodes of the Cayley graph, we verify that performance is not sensitive to the specific permutation. We ran 10 independent random layer-to-node mappings with fixed model seeds on all classification tasks across all three LLMs. Table 3 reports mean and standard deviation across the 10 mappings, showing that performance remains stable with minimal variance, confirming that Cayley-Encoder’s gains are robust to the choice of layer-to-node assignment.
Table 3: Shuffled layer permutation results on Cayley graph topology. We report the mean $\pm$ std over 10 random layer permutations.
Base Model Method Banking77 Emotion MTOP Domain MTOP Intent Poem Sentiment
Pythia-410m Cayley-Encoder (GIN) $89.55\pm 0.34$ $74.91\pm 0.78$ $98.67\pm 0.06$ $94.55\pm 0.39$ $71.26\pm 1.83$
Cayley-Encoder (GCN) $90.07\pm 0.31$ $69.76\pm 1.53$ $98.57\pm 0.09$ $94.31\pm 0.40$ $68.05\pm 0.81$
Gemma2-2B Cayley-Encoder (GIN) $92.35\pm 0.19$ $71.50\pm 2.98$ $99.00\pm 0.08$ $96.22\pm 0.11$ $77.32\pm 2.89$
Cayley-Encoder (GCN) $92.32\pm 0.27$ $73.14\pm 1.46$ $99.14\pm 0.05$ $96.08\pm 0.17$ $82.25\pm 0.54$
Llama3-8B Cayley-Encoder (GIN) $92.35\pm 0.12$ $71.16\pm 0.71$ $98.88\pm 0.08$ $96.42\pm 0.08$ $74.46\pm 1.71$
Cayley-Encoder (GCN) $92.92\pm 0.13$ $72.80\pm 1.30$ $98.90\pm 0.08$ $96.08\pm 0.27$ $79.24\pm 1.31$
In the next subsections, we further examine how the performance of ILSE scales with the LLM size and the number of training samples.
Table 4: Performance comparison of ILSE and baselines across 3 different LLMs on 8 STS tasks. For each task and LLM, we highlight in Bold the best performing setting, in blue the 2nd best and in red the 3rd best. In 20 out of 24 tasks and LLMs, ILSE gets the best score, specifically Cayley-Encoder.
Base Model Section Method STSBenchmark STS12 STS13 STS14 STS15 STS16 BIOSSES SICK-R
Pythia-410m Baselines Last-Layer 39.12 46.96 47.00 41.45 49.32 50.37 67.30 52.55
Best-Layer 53.53 50.62 59.27 51.61 65.59 58.02 74.80 58.26
MLP Last-Layer 25.54 42.14 36.16 33.28 37.75 39.84 59.68 43.23
Weighted 41.81 47.12 49.49 44.21 53.72 52.76 67.79 55.09
DWAtt 49.40 52.51 44.57 47.46 52.51 44.80 45.29 52.43
ILSE Set-Encoder 39.48 52.11 42.68 43.99 48.32 40.48 44.72 44.84
FC-Encoder (GIN) 43.29 44.69 40.21 33.79 54.54 49.56 52.14 46.97
FC-Encoder (GCN) 54.20 50.13 53.00 51.76 63.94 57.11 58.65 57.14
Cayley-Encoder (GIN) 55.84 55.97 60.25 56.65 66.99 58.63 56.59 55.34
Cayley-Encoder (GCN) 49.53 57.90 51.10 52.60 60.51 56.62 39.32 52.15
Gemma2-2B Baselines Last-Layer 36.82 33.70 45.60 37.58 50.22 51.40 58.44 43.01
Best-Layer 52.97 43.02 58.56 52.43 65.49 58.89 72.02 57.24
MLP Last-Layer 40.72 27.36 36.93 31.32 52.38 51.16 54.10 40.19
Weighted 40.52 35.66 45.58 38.14 54.88 54.91 65.27 46.67
DWAtt 53.37 51.86 48.14 55.29 67.73 51.11 57.08 55.36
ILSE Set-Encoder 36.67 43.97 38.97 35.05 50.51 47.34 31.64 42.22
FC-Encoder (GIN) 55.29 46.48 45.79 44.77 67.22 57.28 54.88 54.67
FC-Encoder (GCN) 62.86 61.83 55.43 63.89 74.68 62.96 66.02 60.12
Cayley-Encoder (GIN) 61.62 60.38 63.98 64.59 73.94 62.69 64.07 60.32
Cayley-Encoder (GCN) 55.53 64.36 47.58 57.86 70.80 57.24 61.97 57.52
Llama3-8B Baselines Last-Layer 45.63 32.65 52.66 42.88 58.07 54.32 67.00 51.16
Best-Layer 56.51 47.21 60.09 54.62 66.89 59.04 72.83 56.96
MLP Last-Layer 48.91 34.88 54.80 44.08 61.19 52.55 55.35 46.62
Weighted 45.77 32.86 59.46 54.90 67.30 58.97 52.82 56.91
DWAtt 52.85 51.47 56.26 58.63 66.00 48.86 46.88 51.45
ILSE Set-Encoder 42.31 39.94 52.80 52.00 56.12 39.50 32.96 46.60
FC-Encoder (GIN) 50.87 41.39 34.04 37.66 63.27 53.89 49.62 51.85
FC-Encoder (GCN) 59.33 55.19 53.51 59.88 73.69 57.25 62.81 57.51
Cayley-Encoder (GIN) 63.05 65.31 63.53 70.17 76.96 63.13 55.32 60.74
Cayley-Encoder (GCN) 41.71 47.11 42.48 48.51 52.46 42.09 23.25 44.04
4.4 Effectiveness Across LLM Scales
We evaluate whether the benefits of ILSE persist across LLM sizes. We conduct this evaluation on the classification tasks using the Pythia suite (Biderman et al., 2023): 14M, 70M, 160M, 410M, 1B, 1.4B, and 2.8B parameters. This allows us to analyze how ILSE performance varies with changing model sizes while neutralizing confounding effects due to architectural differences between LLM families (e.g., Pythia vs. LLama).Figure 2 shows performance across Pythia sizes for the classification tasks.
The results demonstrate that ILSE and especially Cayley-Encoder consistently outperforms all baselines across all Pythia sizes. The performance gap between ILSE and Last-Layer and Best-Layer baselines remains consistent even as LLM size increases, with improvements of +20-40% maintained across all Pythia sizes. The smaller LLMs in Pythia suite (14M, 70M and 160M) can reach similar performance as larger ones (1B - 2.8B) using ILSE. For Pythia-14M with MTOP Domain task Cayley-Encoder and FC-Encoder reach 95-96% accuracy just 2-4% lower than ILSE with Pythia-2.8B which reaches 98-99% accuracy. Using Pythia-160M with ILSE we reach to 98% accuracy in MTOP Domain. This trend can be seen also in Banking and MTOP Intent tasks. For Emotion task we see that larger LLMs contribute to increase in accuracy with ILSE, but again ILSE in Pythia-410M reaches 75% accuracy VS 78% in Pythia-1B and 75% in Pythia-2.8B for Emotion task. This suggests ILSE can be effective in low-resource setups with relatively small LLMs.
(a) Banking77
(b) Emotion
(c) MTOP Intent
(d) PoemSentiment
(e) MTOP Domain
Figure 2: LLM size analysis. Performance across the Pythia model suite (14M to 2.8B parameters). ILSE consistently outperforms other baselines across all model sizes.
4.5 Few-Shot Learning
We evaluate the few-shot performance of ILSE by restricting the training data to a range of $1$ to $1024$ samples per label across all classification tasks (for some tasks where training data is limited, we reach the full training data after 128 samples per label).Our results indicate data efficiency. Over 4 tasks including Banking77, MTOP Domain, MTOP Intent and Poem Sentiment, ILSE achieves gains over both the last-layer and best-layer baselines with as few as $8$ samples per label. In Banking77 task using just $32$ samples per label makes Cayley-Encoder and FC-Encoder to surpass all other baselines including DWAtt being trained with full amount of training data. For Poem Sentiment task using 64 samples per label enables ILSE to outperform all other baselines being trained with more data. For Emotion task, starting with $32$ samples per label ILSE outperforms Last-Layer and Best-Layer baselines and for larger training set sizes ILSE outperforms all other baselines per training set size. Emotion is the only task that exhibits consistent performance improvements as the training set size grows for both ILSE and baselines. The other tasks reach plateau between $128$ to $512$ samples per label. The overall trend suggests that ILSE is effective in low-resource settings. This makes it particularly valuable for practical applications where labeled data is scarce (Figure 3).
(a) Banking77
(b) Emotion
(c) MTOP Intent
(d) PoemSentiment
(e) MTOP Domain
Figure 3: Few-Shot Learning Analysis. Performance across 1-1024 samples per label using Pythia-410M. ILSE outperforms baselines with 32 samples per label across all tasks.
4.6 Comparison with LoRA
Although ILSE is designed as a lightweight post-hoc module that is complementary to fine-tuning approaches and can be combined with methods such as LoRA (Hu et al., 2022) or full fine-tuning, we compare it directly against LoRA to demonstrate its standalone effectiveness. LoRA inserts trainable low-rank matrices into the LLM’s attention layers, adapting the model’s internal computation while keeping the original weights frozen. To keep the comparison fair in number of parameters, we capped the LoRa rank to 3. Tables 5 and 6 present the results on classification and STS tasks, respectively. On classification, ILSE outperforms LoRA in 13 out of 15 configurations of task and LLM combination. On STS tasks, Cayley-Encoder outperforms LoRA in 7 out of 8 tasks across all three LLMs. These results show the strength of structured inter-layer representations aggregation.
Table 5: ILSE vs LoRA accuracy on classification tasks. For each encoder, we report the best result between GIN and GCN variants. For each task and LLM, we highlight in Bold the best performing setting.
Base Model Method Banking77 Emotion MTOPDomain MTOPIntent PoemSentiment
Pythia-410m Set-Encoder 84.23 47.89 97.59 92.21 73.37
FC-Encoder 90.65 75.61 98.68 95.04 75.77
Cayley-Encoder 89.43 73.83 98.77 94.72 70.87
LoRA 79.35 91.06 96.64 88.25 70.96
Gemma2-2B Set-Encoder 90.03 78.77 98.92 94.37 78.56
FC-Encoder 92.39 79.90 99.07 95.34 77.12
Cayley-Encoder 92.58 69.60 99.16 96.43 83.27
LoRA 56.77 70.89 89.65 79.64 61.92
Llama3-8B Set-Encoder 87.62 71.04 98.77 95.43 77.02
FC-Encoder 92.38 71.64 98.99 96.19 77.98
Cayley-Encoder 92.85 73.43 99.03 96.46 79.04
LoRA 90.82 91.98 98.59 94.44 76.15
Table 6: ILSE vs LoRA spearman correlation ($\times 100$) on STS tasks. For each encoder, we report the best result between GIN and GCN variants of STSBenchmark. For each task and LLM, we highlight in Bold the best performing setting.
Base Model Method STSBenchmark STS12 STS13 STS14 STS15 STS16 BIOSSES SICK-R
Pythia-410m Set-Encoder $41.0$ $\mathbf{56.0}$ $43.7$ $44.2$ $44.0$ $36.0$ $47.0$ $42.6$
FC-Encoder $54.2$ $50.1$ $53.0$ $51.8$ $63.9$ $57.1$ $58.6$ $\mathbf{57.1}$
Cayley-Encoder $\mathbf{55.8}$ $56.0$ $\mathbf{60.2}$ $\mathbf{56.7}$ $\mathbf{67.0}$ $\mathbf{58.6}$ $56.6$ $55.3$
LoRA $54.1$ $44.6$ $51.7$ $52.3$ $64.8$ $45.8$ $\mathbf{64.3}$ $52.0$
Gemma2-2B Set-Encoder $36.7$ $44.0$ $39.0$ $35.1$ $50.5$ $47.3$ $31.6$ $42.2$
FC-Encoder $\mathbf{62.9}$ $\mathbf{61.8}$ $55.4$ $63.9$ $\mathbf{74.7}$ $\mathbf{63.0}$ $66.0$ $60.1$
Cayley-Encoder $61.6$ $60.4$ $\mathbf{64.0}$ $\mathbf{64.6}$ $73.9$ $62.7$ $64.1$ $\mathbf{60.3}$
LoRA $53.6$ $44.6$ $49.1$ $50.8$ $63.5$ $46.7$ $\mathbf{66.4}$ $52.2$
Llama3-8B Set-Encoder $42.5$ $42.0$ $50.6$ $49.1$ $50.0$ $41.2$ $34.3$ $44.7$
FC-Encoder $59.3$ $55.2$ $53.5$ $59.9$ $73.7$ $57.2$ $62.8$ $57.5$
Cayley-Encoder $\mathbf{63.1}$ $\mathbf{65.3}$ $\mathbf{63.5}$ $\mathbf{70.2}$ $\mathbf{77.0}$ $\mathbf{63.1}$ $55.3$ $\mathbf{60.7}$
LoRA $62.8$ $47.8$ $59.0$ $61.0$ $70.6$ $52.1$ $\mathbf{64.1}$ $59.1$
4.7 Token-Level Inter-Layer Aggregation
In all experiments above, each layer is represented by a single vector obtained via mean-pooling over tokens. While this is effective, it discards fine-grained token-level information. To investigate it, we construct a graph where each node corresponds to a token at a specific layer, resulting in $L\times T$ nodes per input, where $L$ is the number of layers of the base LLM and $T$ is the number of tokens in the input sequence. We apply Cayley-Encoder and FC-Encoder over this graph using Pythia-14M on three classification tasks . Table  7 shows substantial improvements over the mean-pooled variant. On Banking77, accuracy increases from 70–77% to 81–87%, on Emotion from 40–47% to 73–87%, and on Poem Sentiment from 48–55% to 50–58%. However, this comes at a significant computational cost. The number of graph nodes grows from $L$ to $L\times T$, substantially increasing memory and runtime requirements. This setting also highlights the scalability advantage of Cayley-Encoder over FC-Encoder. With $L\times T$ nodes, the fully-connected graph becomes large, with peak memory usage of 828–1516 MB compared to only 77–121 MB for Cayley-Encoder. This demonstrates that the sparse $O(n)$ edge complexity of the Cayley topology is highly beneficial when scaling to larger graphs in compare to the $O(n^{2})$ cost of FC-Encoder.
Table 7: Comparison of pooled-layer vs. all-tokens ILSE methods on Pythia-14m (L=7L{=}7 layers, D=128D{=}128). The original ILSE aggregated tokens via mean pooling per layer resulting, $L$ nodes; All-Tokens use every (layer, token) pair as a node ($L\times T$ nodes). All-tokens results show mean $\pm$ std over 3 seeds. Node ranges are estimated from 10 random samples per task. Peak memory is averaged across seeds and tasks. Cayley-Encoder shows lower peak memory - about 10 fold change.
Base Model Aggregation Method Banking77 Emotion PoemSentiment Nodes Peak Mem (MB)
Pythia-14m Mean-Pooled ($L$ nodes) Cayley-Encoder (GCN) $69.97$ $39.78$ $47.69$ 77
FC-Encoder (GCN) $77.12$ $47.46$ $54.81$ 77
Cayley-Encoder (GIN) $71.00$ $45.97$ $48.27$ 77
FC-Encoder (GIN) $75.54$ $47.30$ $36.73$ 77
All-Tokens ($L\times T$ nodes) Cayley-Encoder (GCN) $85.41\pm 0.47$ $87.88\pm 0.95$ $55.03\pm 0.06$ 2828175175 $121$
FC-Encoder (GCN) $81.74\pm 0.57$ $73.84\pm 1.38$ $50.32\pm 9.47$ 2828175175 1,5161{,}516
Cayley-Encoder (GIN) $7.46\pm 0.42 88$$8.26\pm 0.11$ $58.01\pm 4.69$ 2828175175 $77$
FC-Encoder (GIN) $81.21\pm 0.76$ $74.41\pm 0.93$ $58.17\pm 3.44$ 2828175175 $828$
5 Discussion and Future Work
Our results show that all three ILSE topologies consistently outperform the baselines, demonstrating the value of structured layer aggregation. Set-Encoder proves that even simple permutation-invariant pooling benefits from using all layers. FC-Encoder shows that dense inter-layer communication yields further gains. Cayley-Encoder achieves the overall best performance despite its topology carrying no task-specific signal. This suggests that sparse, regular connectivity regularizes the aggregation, allowing generalization while enabling efficient information flow between all layers. The consistent gains across all three structures indicate that utilizing information from all layers is as important as the choice of aggregation method.
As future work, several promising directions emerge. While we evaluated models up to 8B parameters, future work could investigate whether ILSE holds for larger architectures (e.g., Llama-3-70B) or in generative contexts, such as mitigating hallucinations through enhanced multi-layer representation mixing. Additionally, we can explore alternative mappings from layers to nodes in cayley topology, such as learned permutations, or task-specific mappings for particular downstream tasks.
Finally, ILSE currently utilizes the structures as a task-specific fusion layer for frozen models. A promising direction is the integration of ILSE directly into the transformer backbone during pre-training. Such a ”geometry-aware” architecture could potentially yield more robust base models with improved global information flow.
6 Conclusion
We introduced ILSE, a structured framework for combining representations from all internal layers of a pre-trained LLM into a single downstream representation. Across a broad set of classification and semantic similarity tasks, ILSE consistently outperformed baselines by large margins, including in few-shot settings and across diverse LLM sizes. These gains were achieved with only a very small number of additional trainable parameters relative to the base model size. Among the variants we studied, the geometric Cayley-Encoder achieved the strongest overall performance. This suggests that sparse, regular, and bottleneck-free inter-layer connectivity provides an effective inductive bias for aggregating layer-wise information. Overall, ILSE offers a simple, efficient, and mathematically grounded approach for improving frozen LLM representations, and suggests that the internal depth of LLMs contains complementary task-relevant signals that can be more effectively exploited through structured aggregation.
7 Acknowledgement
This study was supported in part by a fellowship from the Edmond J. Safra Center for Bioinformatics at Tel-Aviv University.
Appendix A Classification Results with Delta Calculations
We include here an extended version of Table 1, which includes the delta percentages between ILSE and the baselines for each LLM. In classification all ILSE methods improve over the baselines, except for one case.
Table 8: Performance comparison of ILSE VS baselines on classification tasks with delta improvements. Bold: best per column, blue: 2nd best, red: 3rd best. Deltas: green = improvement, red = degradation.
Base Model Section Method Banking77 Emotion MTOPDomain MTOPIntent PoemSentiment Avg
Pythia-410m Baselines Last Layer 61.17 33.48 80.88 66.97 42.40 56.98
Best Single Layer 66.67 35.02 83.78 71.18 45.67 60.47
MLP Last Layer 83.84 33.99 96.97 83.75 75.00 74.71
MLP Best Layer 41.93 25.41 87.25 70.79 53.94 55.86
Weighted 58.50 28.26 79.79 61.68 42.60 54.16
DWATT 83.23 58.60 98.03 91.66 70.87 80.48
Structural Fusion Set-Encoder 84.23 47.89 97.59 92.21 73.37 79.06
FC-Encoder (GIN) 90.10 73.36 98.68 94.32 69.90 85.27
FC-Encoder (GCN) 90.65 75.61 98.65 95.04 75.77 87.14
Cayley-Encoder (GIN) 89.12 73.83 98.77 94.72 70.87 85.46
Cayley-Encoder (GCN) 89.43 66.40 98.67 94.19 69.13 83.56
Gemma2-2B Baselines Last Layer 62.16 26.89 80.79 68.02 35.67 54.71
Best Single Layer 72.31 31.27 87.09 75.36 42.79 61.76
MLP Last Layer 87.47 59.05 98.37 92.73 71.63 81.85
MLP Best Layer 59.85 30.26 84.24 73.69 46.83 58.97
Weighted 65.40 29.94 85.17 73.20 39.62 58.67
DWATT 88.82 61.22 98.39 90.93 67.79 81.43
Structural Fusion Set-Encoder 90.03 78.77 98.92 94.37 78.56 88.13
FC-Encoder (GIN) 92.39 79.55 99.07 94.47 74.62 88.02
FC-Encoder (GCN) 91.71 79.90 98.87 95.34 77.12 88.59
Cayley-Encoder (GIN) 92.58 68.38 98.88 96.43 76.44 86.54
Cayley-Encoder (GCN) 91.86 69.60 99.16 95.97 83.27 87.97
Llama3-8B Baselines Last Layer 68.25 34.23 84.42 73.39 40.96 60.25
Best Single Layer 71.93 38.42 89.01 78.17 47.02 64.91
MLP Last Layer 86.70 67.67 98.58 92.09 75.00 84.01
MLP Best Layer 49.59 21.17 47.64 60.04 35.67 42.82
Weighted 66.63 27.94 83.85 71.78 37.79 57.60
DWATT 90.04 66.55 97.97 92.41 75.00 84.40
Structural Fusion Set-Encoder 87.62 71.04 98.77 95.43 77.02 85.98
FC-Encoder (GIN) 92.10 71.64 98.77 95.65 75.96 86.83
FC-Encoder (GCN) 92.38 71.03 98.99 96.19 77.98 87.31
Cayley-Encoder (GIN) 92.46 71.58 98.98 96.46 76.54 87.20
Cayley-Encoder (GCN) 92.85 73.43 99.03 95.90 79.04 88.05
$\Delta$ vs Last Layer Set-Encoder +23.43 +34.36 +16.40 +24.55 +36.63 +27.08
FC-Encoder (GIN) +27.67 +43.32 +16.81 +25.36 +33.81 +29.39
FC-Encoder (GCN) +27.72 +43.98 +16.80 +26.06 +37.28 +30.37
Cayley-Encoder (GIN) +27.53 +39.73 +16.84 +26.41 +34.94 +29.09
Cayley-Encoder (GCN) +27.52 +38.27 +16.92 +25.90 +37.47 +29.22
$\Delta$ vs Best Single Layer Set-Encoder +16.99 +30.99 +11.80 +19.10 +31.15 +22.01
FC-Encoder (GIN) +21.23 +39.94 +12.22 +19.91 +28.33 +24.33
FC-Encoder (GCN) +21.27 +40.60 +12.21 +20.62 +31.79 +25.30
Cayley-Encoder (GIN) +21.08 +36.35 +12.25 +20.97 +29.46 +24.02
Cayley-Encoder (GCN) +21.07 +34.90 +12.33 +20.45 +31.99 +24.15
$\Delta$ vs MLP Last Layer Set-Encoder +1.29 +12.33 +0.46 +4.48 +2.44 +4.20
FC-Encoder (GIN) +5.53 +21.28 +0.87 +5.29 -0.38 +6.52
FC-Encoder (GCN) +5.58 +21.94 +0.86 +6.00 +3.08 +7.49
Cayley-Encoder (GIN) +5.38 +17.69 +0.90 +6.35 +0.74 +6.21
Cayley-Encoder (GCN) +5.37 +16.24 +0.98 +5.83 +3.27 +6.34
$\Delta$ vs DWATT Set-Encoder -0.07 +3.78 +0.30 +2.34 +5.10 +2.29
FC-Encoder (GIN) +4.17 +12.73 +0.71 +3.15 +2.28 +4.61
FC-Encoder (GCN) +4.21 +13.39 +0.71 +3.86 +5.74 +5.58
Cayley-Encoder (GIN) +4.02 +9.14 +0.75 +4.20 +3.40 +4.30
Cayley-Encoder (GCN) +4.01 +7.68 +0.82 +3.69 +5.93 +4.43
$\Delta$ Cayley vs FC Cayley-Encoder (GIN) -0.14 -3.59 +0.03 +1.05 +1.12 -0.30
Cayley-Encoder (GCN) -0.20 -5.70 +0.12 -0.17 +0.19 -1.15
$\Delta$ Avg Cayley vs FC Cayley - FC -0.17 -4.65 +0.08 +0.44 +0.66 -0.73
Appendix B STS Results with Delta Calculations
We include here an extended version of Table 2, which includes the delta percentages between ILSE and the baselines for each LLM. In STS Cayley-Encoder shows dominance among ILSE methods.
Table 9: Performance comparison of ILSE VS baselines on STS tasks with delta improvements. Bold: best per column, blue: 2nd best, red: 3rd best. Deltas: green = improvement, red = degradation.
Base Model Section Method STSBenchmark STS12 STS13 STS14 STS15 STS16 BIOSSES SICK-R Avg
Pythia-410m Baselines Last-Layer 39.12 46.96 47.00 41.45 49.32 50.37 67.30 52.55 49.26
Best-Layer 53.53 50.62 59.27 51.61 65.59 58.02 74.80 58.26 58.96
MLP Last-Layer 25.54 42.14 36.16 33.28 37.75 39.84 59.68 43.23 39.70
Weighted 41.81 47.12 49.49 44.21 53.72 52.76 67.79 55.09 51.50
DWAtt 49.40 52.51 44.57 47.46 52.51 44.80 45.29 52.43 48.62
ILSE Set-Encoder 39.48 52.11 42.68 43.99 48.32 40.48 44.72 44.84 44.58
FC-Encoder (GIN) 43.29 44.69 40.21 33.79 54.54 49.56 52.14 46.97 45.65
FC-Encoder (GCN) 54.20 50.13 53.00 51.76 63.94 57.11 58.65 57.14 55.74
Cayley-Encoder (GIN) 55.84 55.97 60.25 56.65 66.99 58.63 56.59 55.34 58.28
Cayley-Encoder (GCN) 49.53 57.90 51.10 52.60 60.51 56.62 39.32 52.15 52.47
Gemma2-2B Baselines Last-Layer 36.82 33.70 45.60 37.58 50.22 51.40 58.44 43.01 44.60
Best-Layer 52.97 43.02 58.56 52.43 65.49 58.89 72.02 57.24 57.58
MLP Last-Layer 40.72 27.36 36.93 31.32 52.38 51.16 54.10 40.19 41.77
Weighted 40.52 35.66 45.58 38.14 54.88 54.91 65.27 46.67 47.70
DWAtt 53.37 51.86 48.14 55.29 67.73 51.11 57.08 55.36 54.99
ILSE Set-Encoder 36.67 43.97 38.97 35.05 50.51 47.34 31.64 42.22 40.80
FC-Encoder (GIN) 55.29 46.48 45.79 44.77 67.22 57.28 54.88 54.67 53.30
FC-Encoder (GCN) 62.86 61.83 55.43 63.89 74.68 62.96 66.02 60.12 63.47
Cayley-Encoder (GIN) 61.62 60.38 63.98 64.59 73.94 62.69 64.07 60.32 63.95
Cayley-Encoder (GCN) 55.53 64.36 47.58 57.86 70.80 57.24 61.97 57.52 59.11
Llama3-8B Baselines Last-Layer 45.63 32.65 52.66 42.88 58.07 54.32 67.00 51.16 50.54
Best-Layer 56.51 47.21 60.09 54.62 66.89 59.04 72.83 56.96 59.27
MLP Last-Layer 48.91 34.88 54.80 44.08 61.19 52.55 55.35 46.62 49.80
Weighted 45.77 32.86 59.46 54.90 67.30 58.97 52.82 56.91 53.62
DWAtt 52.85 51.47 56.26 58.63 66.00 48.86 46.88 51.45 54.05
ILSE Set-Encoder 42.31 39.94 52.80 52.00 56.12 39.50 32.96 46.60 45.28
FC-Encoder (GIN) 50.87 41.39 34.04 37.66 63.27 53.89 49.62 51.85 47.82
FC-Encoder (GCN) 59.33 55.19 53.51 59.88 73.69 57.25 62.81 57.51 59.90
Cayley-Encoder (GIN) 63.05 65.31 63.53 70.17 76.96 63.13 55.32 60.74 64.77
Cayley-Encoder (GCN) 41.71 47.11 42.48 48.51 52.46 42.09 23.25 44.04 42.71
$\Delta$ vs Last-Layer Set-Encoder -1.04 +7.57 -3.60 +3.04 -0.88 -9.59 -27.81 -4.36 -4.58
FC-Encoder (GIN) +9.29 +6.42 -8.41 -1.90 +9.14 +1.54 -12.04 +2.25 +0.79
FC-Encoder (GCN) +18.27 +17.95 +5.56 +17.87 +18.24 +7.07 -1.75 +9.35 +11.57
Cayley-Encoder (GIN) +19.65 +22.78 +14.17 +23.17 +20.09 +9.45 -5.59 +9.89 +14.20
Cayley-Encoder (GCN) +8.40 +18.69 -1.37 +12.35 +8.72 -0.05 -22.73 +2.33 +3.29
$\Delta$ vs Best-Layer Set-Encoder -14.85 -1.61 -14.49 -9.21 -14.34 -16.21 -36.77 -12.93 -15.05
FC-Encoder (GIN) -4.52 -2.76 -19.29 -14.15 -4.32 -5.08 -21.00 -6.32 -9.68
FC-Encoder (GCN) +4.46 +8.77 -5.33 +5.62 +4.78 +0.45 -10.72 +0.77 +1.10
Cayley-Encoder (GIN) +5.84 +13.60 +3.28 +10.92 +6.64 +2.83 -14.56 +1.31 +3.73
Cayley-Encoder (GCN) -5.41 +9.51 -12.25 +0.10 -4.73 -6.67 -31.70 -6.25 -7.18
$\Delta$ vs MLP Last-Layer Set-Encoder +1.10 +10.55 +2.18 +7.45 +1.21 -5.41 -19.93 +1.21 -0.21
FC-Encoder (GIN) +11.42 +9.39 -2.62 +2.52 +11.23 +5.73 -4.16 +7.82 +5.17
FC-Encoder (GCN) +20.40 +20.92 +11.35 +22.28 +20.33 +11.25 +6.12 +14.91 +15.95
Cayley-Encoder (GIN) +21.78 +25.76 +19.95 +27.58 +22.19 +13.63 +2.28 +15.45 +18.58
Cayley-Encoder (GCN) +10.53 +21.66 +4.42 +16.77 +10.81 +4.13 -14.86 +7.89 +7.67
$\Delta$ vs DWAtt Set-Encoder -12.39 -6.61 -4.84 -10.11 -10.43 -5.82 -13.31 -8.53 -9.00
FC-Encoder (GIN) -2.06 -7.76 -9.64 -15.05 -0.41 +5.32 +2.46 -1.92 -3.63
FC-Encoder (GCN) +6.92 +3.77 +4.32 +4.72 +8.69 +10.85 +12.74 +5.18 +7.15
Cayley-Encoder (GIN) +8.30 +8.60 +12.93 +10.01 +10.55 +13.22 +8.91 +5.72 +9.78
Cayley-Encoder (GCN) -2.95 +4.51 -2.60 -0.80 -0.83 +3.72 -8.24 -1.84 -1.13
$\Delta$ Cayley vs FC Cayley-Encoder (GIN) +10.36 +16.37 +22.57 +25.06 +10.96 +7.91 +6.45 +7.64 +13.41
Cayley-Encoder (GCN) -9.87 +0.74 -6.93 -5.52 -9.52 -7.12 -20.98 -7.02 -8.28
Appendix C Hardware Details
We worked on HPC cluster with access to the following GPUs: A6000, A100, V100 and geforce rtx 3090.
Appendix D Train Time
We couldn’t compare training time properly because our experiments ran on different GPUs, depeneding on the cluster availability. However we decided to add this table for rough estimates of training times among the different methods.
Table 10: Average training time for ISLE and baselines across all LLMs. Times are approximate due to GPU hardware variability (A6000, A100, V100, RTX 3090).
Section Method Bank77 Emot. MTOP-D MTOP-I Poem STS-B Avg
Baselines Weighted 1.0m 1.3m 1.2m 1.3m 5.0s 1.1h 11.5m
DWATT 11.4m 1.1h 26.3m 57.3m 3.0m 57.8m 36.5m
MLP 11.0s 19.1s 19.2s 18.3s 1.4s 2.8m 39.3s
ILSE Set-Encoder 42.6s 2.5m 1.7m 2.6m 5.7s 1.5h 16.7m
FC-Encoder (GIN) 2.6m 4.5m 4.3m 4.8m 17.9s 52.8m 11.5m
FC-Encoder (GCN) 3.0m 5.4m 4.8m 4.4m 17.6s 55.2m 12.2m
Cayley-Encoder (GIN) 3.4m 6.1m 7.4m 5.8m 18.4s 3.2h 35.4m
Cayley-Encoder (GCN) 4.2m 7.4m 6.8m 6.4m 26.1s 1.7h 21.6m